# Why Do Open-Source LLMs Struggle with Data Analysis? A Systematic Empirical Study

### v1

@@ 3.1 Data Collection and Curation
We evaluate a series of open-source model variants, including Qwen2.5-7B-Instruct, Qwen2.5-7B-Coder, Qwen2.5-7B-Instruct-1M, and R1-Distill-Qwen-7B, alongside well-established open-source models like GPT-4o, DeepSeek-v3111We use gpt-4o-2024-08-06 and deepseek-v3-0324., and DeepSeek-R1. Following prior work (Hu et al., 2024; Zeng et al., 2025), we adopt ReAct (Yao et al., 2023) to iteratively alternate between planning, code generation, and execution for multi-turn data analysis. Model training is conducted based on Qwen2.5-Instruct, utilizing LoRA Hu et al. (2022) for efficient fine-tuning (Details on hyperparameters are provided in Appendix A).
@@ 4.3.1 Interaction Turns
(ii) Quality Over Quantity.
@@
(iii) A balanced selection strategy is necessary.
@@ 6 Related Work
Data science is an interdisciplinary field that focuses on extracting valuable insights from various data sources, with data analysis serving as a crucial component. Recently, researchers have proposed several specialized LLMs to enhance automated data analysis capabilities. Data Copilot Zhang et al. (2023) is a code-centric data analysis agent that efficiently handles massive data processing and visualization through pre-designed interfaces, specifically tailored to automate financial data analysis tasks while reducing errors and improving efficiency. Data Interpreter Hong et al. (2024) takes a different path by using hierarchical dependency graphs to represent workflows, enabling automatic task breakdown and code improvements. AutoKaggle Li et al. (2024) coordinates multiple specialized agents (Planner, Developer, and Reviewer) to handle end-to-end data analysis tasks, while keeping room for human input when needed. To evaluate the ability of these agents, researchers have developed several comprehensive benchmarks, such as InfiAgent-DABench Hu et al. (2024), DSBench Jing et al. (2024), DA-code Huang et al. (2024), and DataSciBench Zhang et al. (2025) for assessing the effectiveness of LLM-based data analysis solutions.
@@ 7 Conclusion and Future Work
While our experiments provide valuable insights into the performance of data efficiency in the context of data analysis, several limitations should be acknowledged. Our work does not incorporate experiments on a broader range of foundation models, such as LLaMA Dubey et al. (2024) and other emerging alternatives, which may exhibit distinct characteristics and behaviors. Second, due to GPU limitations, we were unable to extend our experiments to larger-scale datasets that could better reflect the diversity and complexity of real-world data analysis scenarios. Third, our evaluation is restricted to models of specific parameter sizes, leaving the behavior of significantly larger models (e.g., 72B) or smaller models(e.g., 2B) unexplored. Finally, the generalizability of our findings to other tasks, domains and application contexts remains an open question and warrants further investigation.
@@
Our models are fine-tuned using the LLaMA Factory framework (Zheng et al., 2024). Experiments are conducted on 4 NVIDIA A800 80G GPUs, leveraging the DeepSpeed ZeRO-3 optimization for efficient distributed training. We adopt a per-device batch size of 1, with gradient accumulation steps set to 4. Training is performed using bfloat16 precision, and a warmup ratio of 0.02. We employ a cosine learning rate scheduler with an initial learning rate of $1\times 10^{-5}$. All our fine-tuned models are evaluated using vLLM Kwon et al. (2023) with the temperature set to 0.

### v2

@@ 3.1 Data Collection and Curation
We evaluate a series of open-source model variants, including Qwen2.5-7B-Instruct, Qwen2.5-7B-Coder, Qwen2.5-7B-Instruct-1M, and R1-Distill-Qwen-7B, alongside strong baselines like GPT-4o, DeepSeek-v3111We use gpt-4o-2024-08-06 and deepseek-v3-0324., and DeepSeek-R1. Following prior work (Hu et al., 2024; Zeng et al., 2025), we adopt ReAct (Yao et al., 2023) to iteratively alternate between planning, code generation, and execution for multi-turn data analysis. During the subsequent experimental analysis, we adopt two approaches: prompt-based analysis and fine-tuning using Low-Rank Adaptation (LoRA) Hu et al. (2022). Model training is based on Qwen2.5-Instruct, with hyperparameters, experimental configurations, and their rationale detailed in Appendix A.
@@ 4.3.1 Interaction Turns
(ii) Data quality outweighs quantity.
@@
(iii) Mixed strategies demand thoughtful optimization.
@@ 6 Related Work
Data science is an interdisciplinary field that focuses on extracting valuable insights from various data sources, with data analysis serving as a crucial component. Recently, researchers have proposed several specialized LLMs to enhance automated data analysis capabilities Sun et al. (2024); Li et al. (2025). Data Copilot Zhang et al. (2023) is a code-centric data analysis agent that efficiently handles massive data processing and visualization through pre-designed interfaces, specifically tailored to automate financial data analysis tasks while reducing errors and improving efficiency. Data Interpreter Hong et al. (2024) takes a different path by using hierarchical dependency graphs to represent workflows, enabling automatic task breakdown and code improvements. AutoKaggle Li et al. (2024) coordinates multiple specialized agents (Planner, Developer, and Reviewer) to handle end-to-end data analysis tasks, while keeping room for human input when needed. To evaluate the ability of these agents, researchers have developed several comprehensive benchmarks, such as InfiAgent-DABench Hu et al. (2024), DSBench Jing et al. (2024), DA-code Huang et al. (2024), and DataSciBench Zhang et al. (2025) for assessing the effectiveness of LLM-based data analysis solutions.
@@ 7 Conclusion and Future Work
While our experiments provide valuable insights into the performance of data efficiency in the context of data analysis, several limitations should be acknowledged. Our work does not incorporate experiments on a broader range of foundation models, such as LLaMA Dubey et al. (2024) and other emerging alternatives, which may exhibit distinct characteristics and behaviors. Second, due to GPU limitations, we were unable to extend our experiments to larger-scale datasets that could better reflect the diversity and complexity of real-world data analysis scenarios. Third, our evaluation is restricted to models of specific parameter sizes, leaving the behavior of significantly larger models (e.g., 72B) or smaller models(e.g., 2B) unexplored. Data analysis, as a domain, presents unique challenges compared to tasks like mathematical reasoning. These challenges include multi-turn interactions, sophisticated code reasoning, and external environment engagement, which demand data of higher diversity and complexity. Although we have not yet discovered domain-specific data capable of dramatically improving model performance—as has been the case in mathematical reasoning—our curated dataset provides valuable prior knowledge and new perspectives. Addressing these limitations in future work, through richer datasets and broader model testing, will be essential to advancing the scalability and adaptability of LLMs in data analysis and complex application domains.
@@
Task-Specific Details. We adopt different strategies based on the characteristics of each task. For Data Comprehension and Code Capability, we employ prompt-based evaluation methods as these tasks can be effectively assessed without additional training. In the case of Data Comprehension, we observe that when provided with file data, the model often redundantly attempts to reload the table using Python, leading to context overflow and reduced task accuracy. Moreover, fine-tuning may not reliably enhance the model’s ability to comprehend data and may inadvertently reinforce these redundant behaviors, making prompt-based evaluation a more suitable choice. For Code Capability, the current selection of models already supports effective evaluation of coding abilities, making fine-tuning unnecessary. For Strategic Planning, we use LoRA Hu et al. (2022) fine-tuning since this task requires the model to learn complex reasoning patterns and long-term dependencies through targeted training. To ensure fair comparisons across experiments, dataset sizes are controlled to be equal within the same task setting. However, some unavoidable differences exist across tasks due to varying data requirements and availability.
Training and Evaluation Details. For fine-tuning, we employ the LLaMA Factory framework (Zheng et al., 2024). Experiments are conducted on 4 NVIDIA A800 80G GPUs, leveraging the DeepSpeed ZeRO-3 optimization for efficient distributed training. We adopt a per-device batch size of 1, with gradient accumulation steps set to 4. Training is performed using bfloat16 precision, and a warmup ratio of 0.02. We employ a cosine learning rate scheduler with an initial learning rate of $1\times 10^{-5}$. All our fine-tuned models are evaluated using vLLM Kwon et al. (2023) with the temperature set to 0.